OpenStudy (anonymous):

Fill in the missing term: (x - 8)2 = x2 + ____ + 64. A. 4x B. 8x C. -16x D. 32x

OpenStudy (anonymous):

2x-16=2x+?+64

OpenStudy (anonymous):

-?=80

OpenStudy (anonymous):

idk if u mean 2x or x2 or x^2 .....cannot answer

OpenStudy (anonymous):

they are ^2

OpenStudy (anonymous):

(x - 8)^2 = x^2 + ____ + 64.

OpenStudy (anonymous):